OpenStudy (anonymous):

What's the lcm for 81,7, 9, and 3?

OpenStudy (mathhelp346):

...15309?

OpenStudy (mathhelp346):

that's the lowest one i could find

OpenStudy (zarkon):

567

myininaya (myininaya):

lcm of 3 and 9 is 9 lcm of 9 and 7 is 63 lcm of 63 and 81 is (9*9*7)

myininaya (myininaya):

i always get scared when you are fixing to say something zarkon

OpenStudy (zarkon):

I thought of it like this... 3 and 9 divide 81 so 81 is the lcm of those three numbers and 81 and 7 are relatively prime so lcm(81,7)=81*7

OpenStudy (zarkon):

;)

myininaya (myininaya):

how do i channel you zarkon when i need your brain?

OpenStudy (zarkon):

Just think VOLTRON!!

myininaya (myininaya):

what is this voltron is this your imaginary world or some world someone else developed for you

OpenStudy (zarkon):

http://en.wikipedia.org/wiki/Voltron defender of the universe...I am in a constant battle with voltron...some day I will destroy it muhahahahahah....sorry

myininaya (myininaya):

you are hilarious

OpenStudy (zarkon):

thank you for saying hilarious and not crazy :)

myininaya (myininaya):

no you aren't crazy imaginary worlds are fine we believe in the imaginary numbers so why not

myininaya (myininaya):

its logical to assume that an imaginary world exists

OpenStudy (zarkon):

good point...so we are all crazy

myininaya (myininaya):

exactly

myininaya (myininaya):

we are all complex numbers in this complex set

myininaya (myininaya):

are lets not say say we are numbers

myininaya (myininaya):

lets say we are complex elements

OpenStudy (anonymous):

what is wrong with \[81=3^4,9=3^2,7=7,3=3\] making lcm as \[3^4\times 7\]?

OpenStudy (zarkon):

That is so complexified

OpenStudy (anonymous):

but it is a hot number

OpenStudy (zarkon):

only crazy people write out prime factorizations.

OpenStudy (zarkon):

crazy I tell you

OpenStudy (anonymous):

then i must truly be nuts because that is the way i always do it

OpenStudy (zarkon):

oh...ok...it's all good

OpenStudy (zarkon):

I'm grading right now so I'm doing all I can to stay in a good mood ;)